Linear differential equation with constant coefficient


d³y/dx³-13dy/dx+12y





1
Expert's answer
2021-01-15T15:05:25-0500

Solution

For solving differential equation

d³y/dx³-13dy/dx+12y =0

we get the characteristic polynomial

s3 -13*s+12=0

In another form

(s-1)*(s2 +s-12)=0 => Roots s1 = 1, s2 = 3, s3 = -4.

The superposition principle for homogeneous equations then gives us that the general solution to the DE is

y(x) = C1*ex+C2*e3x+C3*e-4x , where C1,C2,C3 are arbitrary constants.  

Answer

y(x) = C1*ex+C2*e3x+C3*e-4x